OpenStudy (anonymous):

Simplify and factor the polynomial n(n+7)+12. Show the original polynomial and the factored form describe the same sequence of the nimbers for n=0, 1, 2, 3, & 4 PLEASE SHOW WORK! : )

OpenStudy (anonymous):

n(n+7) +12 =n^2+7n+12 =(n+3)(n+4) Factored form... this is the first part anyway.